# Towards Affordance-Aware Robotic Dexterous Grasping with Human-like Priors

### v3

@@ 4.2 Metrics
Following previous works (Xu et al. 2023; Wan et al. 2023; Wang et al. 2025), each object is randomly rotated and dropped onto the table to enhance the diversity of its initial poses. We report the success rate of grasp $Succ$, Human-likeness Score $HLS$, and Affordance Score $AS$ across all objects and grasp attempts. A grasp is considered successful if the object reaches the target goal within 200 steps in simulator. The Human-likeness Score $HLS$ assesses the anthropomorphic quality of the grasp, which is obtained by prompting the Gemini 2.5 Pro (Comanici et al. 2025) to analyze a visual sequence of the grasp execution. This metric is specifically to rate the resemblance of the dexterous hand’s motion to that of a typical human, yielding a quantitative measure of naturalness. The Affordance Score $AS$, in contrast, evaluates the functional correctness of the grasp by penalizing contact with inappropriate object parts. This metric is calculated using a point cloud of 100 “negative affordance” points sampled from our NAA. Specifically, the score is incremented by one for each fingertip that maintains a distance greater than 2cm from any point in this negative set, thus rewarding functionally sound grasps.
@@ 4.3 Implementation Details
We conduct our experiments in IssacGym (Makoviychuk et al. 2021) simulator. During training, 4096 environments are simulated in parallel on an NVIDIA RTX 4090 GPU. For the network architecture, we use MLP with 4 hidden layers (1024,1024,512,512) for the policy network and value network in the state-based setting, and an additional PointNet+Transformer (Mu et al. 2021) to encode the 3D scene point cloud input in the vision-based setting. Other detailed hyperparameters are shown in our Supp. Mat.
@@ 4.4 Comparison with SOTA Methods
We evaluate AffordDex by first training a state-based policy and then distilling it into a vision-based one. For comparison, we compare AffordDex with several state-of-the-art methods. These include RL and imitation learning algorithms like PPO (Schulman et al. 2017), DAPG (Rajeswaran et al. 2017), and ILAD (Wu, Wang, and Wang 2023). We also compare against methods with advanced learning paradigms such as GSL (Jia et al. 2022) (generalist-specialist), UniDexGrasp++ (Wan et al. 2023) (geometry-aware curriculum learning), UniDexGrasp++ (Wan et al. 2023) which proposes a geometry-aware curriculumn learning and generalist-specialist learning, and DexGrasp Anything (Zhong et al. 2025), diffusion-based dexterous grasp generation models.
Since DexGrasp Anything (Zhong et al. 2025) generates only the final static grasp pose rather than a grasping motion, the $HLS$ is not applicable in this setting. To evaluate grasp robustness, we apply a random external force ranging from 0 to 200N to the object to simulate object’s gravity.
@@ 4.5 Ablation study
NAA. As shown in Fig. 7, a naive approach combining an MLLM (Achiam et al. 2023) with SAM (Kirillov et al. 2023) (denoted as GPT+SAM) proves ineffective for this task. This baseline first uses the MLLM’s coarse localization ability to provide prompts to SAM. However, because MLLMs like GPT-4V (Achiam et al. 2023) excel at image-level understanding but struggle with the fine-grained spatial localization required for segmentation, this process often results in the segmentation of the entire object. In contrast, our NAA module solves this by converting the segmentation task into a simpler classification problem. By first using SAM to generate accurate mask proposals and then using CLIP to select the one with the highest semantic similarity to the negative affordance description, NAA achieves precise segmentation.
@@ 4.6 Extension to Other Grasping Methods
Notably, our proposed modules demonstrate strong generalizability by significantly enhancing other RL-based methods, such as UniDexGrasp++ (Wan et al. 2023). Specifically, the Human Hand Trajectory Imitating (HTI) module markedly improves the naturalness and human-likeness of its generated poses. Simultaneously, the Affordance-aware Residual Learning, guided by negative affordances from our Negative Affordance-aware Segmentation (NAA), substantially boosts the semantic appropriateness of its grasp locations on the object, as shown in Tab. 3.

### v4

@@ 4.2 Metrics
Following previous works (xu2023unidexgrasp; wan2023unidexgrasp++; wang2025unigrasptransformer), each object is randomly rotated and dropped onto the table to enhance the diversity of its initial poses. We report the success rate of grasp $Succ$, Human-likeness Score $HLS$, and Affordance Score $AS$ across all objects and grasp attempts. A grasp is considered successful if the object reaches the target goal within 200 steps in simulator. The Human-likeness Score $HLS$ assesses the anthropomorphic quality of the grasp, which is obtained by prompting the Gemini 2.5 Pro (comanici2025gemini) to analyze a visual sequence of the grasp execution. This metric is specifically to rate the resemblance of the dexterous hand’s motion to that of a typical human, yielding a quantitative measure of naturalness. The Affordance Score $AS$, in contrast, evaluates the functional correctness of the grasp by penalizing contact with inappropriate object parts. This metric is calculated using a point cloud of 100 “negative affordance” points sampled from our NAA. Specifically, the score is incremented by one for each fingertip that maintains a distance greater than 2cm from any point in this negative set, thus rewarding functionally sound grasps.
@@ 4.3 Implementation Details
We conduct our experiments in IssacGym (makoviychuk2021isaac) simulator. During training, 4096 environments are simulated in parallel on an NVIDIA RTX 4090 GPU. For the network architecture, we use MLP with 4 hidden layers (1024,1024,512,512) for the policy network and value network in the state-based setting, and an additional PointNet+Transformer (mu2021maniskill) to encode the 3D scene point cloud input in the vision-based setting. Other detailed hyperparameters are shown in our Supp. Mat.
@@ 4.4 Comparison with SOTA Methods
We evaluate AffordDex by first training a state-based policy and then distilling it into a vision-based one. For comparison, we compare AffordDex with several state-of-the-art methods. These include RL and imitation learning algorithms like PPO (schulman2017proximal), DAPG (rajeswaran2017learning), and ILAD (wu2023learning). We also compare against methods with advanced learning paradigms such as GSL (jia2022improving) (generalist-specialist), UniDexGrasp++ (wan2023unidexgrasp++) (geometry-aware curriculum learning), UniDexGrasp++ (wan2023unidexgrasp++) which proposes a geometry-aware curriculumn learning and generalist-specialist learning, and DexGrasp Anything (zhong2025dexgrasp), diffusion-based dexterous grasp generation models.
Since DexGrasp Anything (zhong2025dexgrasp) generates only the final static grasp pose rather than a grasping motion, the $HLS$ is not applicable in this setting. To evaluate grasp robustness, we apply a random external force ranging from 0 to 200N to the object to simulate object’s gravity.
@@ 4.5 Ablation study
NAA. As shown in Fig. 7, a naive approach combining an MLLM (achiam2023gpt) with SAM (kirillov2023segment) (denoted as GPT+SAM) proves ineffective for this task. This baseline first uses the MLLM’s coarse localization ability to provide prompts to SAM. However, because MLLMs like GPT-4V (achiam2023gpt) excel at image-level understanding but struggle with the fine-grained spatial localization required for segmentation, this process often results in the segmentation of the entire object. In contrast, our NAA module solves this by converting the segmentation task into a simpler classification problem. By first using SAM to generate accurate mask proposals and then using CLIP to select the one with the highest semantic similarity to the negative affordance description, NAA achieves precise segmentation.
@@ 4.6 Extension to Other Grasping Methods
Notably, our proposed modules demonstrate strong generalizability by significantly enhancing other RL-based methods, such as UniDexGrasp++ (wan2023unidexgrasp++). Specifically, the Human Hand Trajectory Imitating (HTI) module markedly improves the naturalness and human-likeness of its generated poses. Simultaneously, the Affordance-aware Residual Learning, guided by negative affordances from our Negative Affordance-aware Segmentation (NAA), substantially boosts the semantic appropriateness of its grasp locations on the object, as shown in Tab. 3.
@@ 5 Conclusion
6 Acknowledgement
This work was supported by Central Guidance for Local Science and Technology Development Fund (ZYYD2025QY19), an Bingtuan Science and Technology Program (2022DB005). This work was supported by DAMO Academy through Academy Research Intern Program.
7 Details about Our Method
Algorithm 1 Overall AffordDex Framework
1:Human hand motion dataset $\mathcal{D}_{H}$, object dataset $\mathcal{D}_{O}$
2:Final vision-based grasping policy $\pi^{S}$
3:Stage 1: Human Hand Trajectory Imitating
4:Pre-train a base policy $\pi^{H}$ on $\mathcal{D}_{H}$ via imitation learning to establish a human motion prior.
5:Stage 2: Affordance-aware Residual Learning and Distillation
6:Generate negative affordance point clouds $\{n_{O}\}_{O\in\mathcal{D}_{O}}$ for all objects using the NAA module.
7:Learn a state-based teacher policy $\pi^{T}$ by fine-tuning $\pi^{H}$ with a residual module, guided by the negative affordances $\{n_{O}\}$.
8:Distill the teacher policy $\pi^{T}$ into a vision-based student policy $\pi^{S}$ via a teacher-student framework (e.g., DAgger).
9:return Final vision-based policy $\pi^{S}$.
8 Details about Baselines
PPO. We use Proximal Policy Optimization (PPO) (schulman2017proximal), a standard on-policy reinforcement learning (RL) algorithm, as the foundation for our RL-based training stages.
DAPG. To leverage expert data, we employ Demo-Augmented Policy Gradient (DAPG) (rajeswaran2017learning). This imitation learning (IL) method accelerates policy training by combining a policy gradient loss with a behavior cloning loss on expert demonstrations. These demonstrations are generated via motion planning.
GSL. We also include Generalist-Specialist Learning (GSL) (jia2022improving), a three-stage training paradigm. It first trains a generalist policy, then fine-tunes specialized policies on difficult task subsets, and finally uses demonstrations from these specialists to train a final, more capable generalist via IL. For fair comparison, our GSL implementation uses PPO for the RL components and DAPG for the final IL stage.
ILAD. To further improve generalization, we compare against ILAD (wu2023learning), an IL method that builds upon DAPG. It introduces an auxiliary objective that forces the policy to learn a geometric object representation from the same motion-planned demonstrations, enhancing its adaptability.
UniDexGrasp and UniDexGrasp++. UniDexGrasp (xu2023unidexgrasp) is a two-stage learning method for dexterous grasping. In the first stage, it trains a state-based teacher policy using Reinforcement Learning. To manage training across numerous objects, it introduces Object Curriculum Learning (OCL), which starts with a single object and gradually incorporates more objects from similar semantic categories. In the second stage, this proficient teacher policy is distilled into a vision-based student policy using DAgger (ross2011reduction), enabling it to operate from point cloud inputs.
Building upon this foundation, UniDexGrasp++ aims to significantly enhance the policy’s generalizability across thousands of object instances with diverse geometries. It argues that curriculum based on semantic categories can be suboptimal and instead proposes two novel, geometry-centric techniques: Geometry-aware Curriculum Learning (GeoCurriculum) and Geometry-aware iterative Generalist-Specialist Learning (GiGSL). GeoCurriculum organizes the training progression based on object geometric features rather than categories, creating a more effective learning path for grasping. GiGSL further refines the policy by iteratively training a generalist model on all objects and specialist models on geometrically challenging subsets. These innovations lead to a more robust universal grasping policy that substantially outperforms its predecessor.
DexGrasp Anything. DexGrasp Anything (zhong2025dexgrasp) is a recent method for generating high-quality, static dexterous grasp poses. It utilizes a diffusion-based generative model to address object diversity and hand complexity. Its core innovation is the direct integration of physical constraints (e.g., collision-freeness, stability) into both the training and sampling phases of the diffusion process. Since it generates only the final static grasp pose rather than a grasping motion, we apply a random external force ranging from 0 to 200N to the object to simulate object’s gravity.
For PPO, DAPG, GSL, ILAD, UniDexGrasp, and UniDexGrasp++, we maintain the same experimental settings as reported in the UniDexGrasp++ paper (wan2023unidexgrasp++) to ensure a fair comparison. For the DexGrasp Anything baseline, we directly utilized the officially released, pretrained model weights provided by the authors.
9 Experiment Details
We use PPO to train Reinforcement learning. We use DAgger-based policy distillation to distill the state-based policy into a vision-based one. We set $w_{f}=1$ in Eq. 1. We follow FGVP and set NMS IoU threshold in NAA to 0.7 (robust in practice).
For UniDexGrasp in Seen Object, we conduct 5 independent runs using fixed random seed of 42, to assess the performance consistency, reported in the table below.
Figure 8: Illustration of the simulation environment.
9.1 Experiment Setup
State Definition. The full state of the teacher policy $S^{T}_{t}=\{R_{t},O_{t},P_{t},N_{t}\}$. The complete object point clouds are assumed to be perfectly accurate. Object states $O_{t}$, including positions, rotations, and velocities, are directly accessible. To accelerate the training process, we sample 1024 points from the object and the hand in the scene point cloud $P_{t}$.
The full state of the student policy $S^{S}_{t}=\{R_{t},P_{t},N_{t}\}$. Partial object point clouds are reconstructed and segmented from depth data captured by five cameras around the table. The hand-object distance is computed using the partial object point cloud in the vision-based setting.
Action Space. The action space consists of the motor commands for the 24 actuators of the dexterous hand. The first 6 actuators control the global position and orientation, while the remaining 18 control the finger joints. We normalize the action range to (1,1)(-1,1).
Camera Setup. Following a similar approach to UniDexGrasp++ (wan2023unidexgrasp++), five RGBD cameras are mounted around the table, as shown in Fig. 8. The cameras are positioned relative to the table center at coordinates $(0.0,0.0,0.55)$, $(0.5,0.0,0.15)$, (0.5,0.0,0.15)(-0.5,0.0,0.15), $(0.0,0.5,0.15)$, (0.0,0.5,0.15)(0.0,-0.5,0.15), with their focal points aligned at $(0,0,0.15)$. In the vision-based setting, the depth images captured by these cameras are fused to generate a scene point cloud, from which the partial point cloud of the object is segmented.
Table 4: Ablation Study on UniDexGrasp dataset (xu2023unidexgrasp) in Seen Object in state-based setting.
Configuration $Succ\uparrow$ $HLS\uparrow$ $AS\downarrow$
$\lambda^{smooth}=0.02$ 89.0 7.9 4
$\lambda^{smooth}=0.05$ 89.2 8.6 4
$\lambda^{smooth}=0.1$ 88.2 8.5 6
$\lambda^{finger}=0.5$ 87.8 8.5 4
$\lambda^{finger}=0.8$ 89.2 8.6 4
$\lambda^{finger}=1.0$ 88.5 8.2 6
Reward Function for Human Hand Trajectory Imitating. The goal of our human hand trajectory imitation reward, $r^{H}$ is to encourage the agent to mirror the articulation and motion smoothness of a reference human trajectory. It is formulated as a weighted sum of two key components: a finger imitation reward and a smoothness reward.
The finger imitation reward $r^{H}_{\text{finger}}$ encourages the agent’s hand to accurately track the reference finger poses. We define this reward based on the squared Euclidean distance between corresponding keypoints on the agent’s hand and the human reference hand:
where $\mathbf{j}_{d,f}$ is the position of the $f$-th keypoint on the dexterous hand, and $\mathbf{j}_{h,f}$ is its corresponding target position from the reference trajectory. The parameters $w_{f}$ and $\lambda_{f}$ are set according to the anatomical group to which keypoint $f$ belongs. Specifically, we group keypoints into two levels:
  • For keypoints $f$ in Level-1 (base joints), we use a stricter decay rate of $\lambda_{f}=50$.
  • For keypoints $f$ in Level-2 (middle joints), we use $\lambda_{f}=40$.
The weights $w_{f}$ are configured to combine these rewards appropriately. This hierarchical parameterization allows the model to prioritize accurate positioning of the more critical base joints.
The smoothness reward $r^{H}_{\text{smooth}}$ is designed to alleviate jerky motions, penalizing the power exerted on each joint, defined as the element-wise product of joint velocities and torques
$r_{t}^{\text{smooth}}=-w_{\text{smooth}}\sum_{i=1}^{n}\left|\tau_{t,i}\cdot\dot{q}_{t,i}\right|,$ (7)
where $w_{\text{smooth}}$ is a positive weighting coefficient, $n$ is the total number of actuated joints, and $\tau_{t,i}$ and $\dot{q}_{t,i}$ are the torque and angular velocity of the $i$-th joint at timestep $t$, respectively. The product $\tau_{t,i}\cdot\dot{q}_{t,i}$ represents the instantaneous power exerted by the actuator of joint $i$. By taking the absolute value, we penalize any high-power action, including both strong acceleration and aggressive braking, thus encouraging the policy to learn energy-efficient and hardware-friendly motions.
The final reward in this stage is:
$r^{H}=\lambda^{\text{smooth}}r_{t}^{\text{smooth}}+\lambda^{\text{finger}}r^{H}_{\text{finger}},$ (8)
where we set $\lambda^{\text{smooth}}=0.05$ and $\lambda^{\text{finger}}=0.8$. We also conduct ablation studies on these weights in Tab. 4.
Reward Function for Affordance-aware Residual Learning. The reward function described in Eq.(4) of the main paper comprises four components: a grasp reward $r_{d}^{T}$, a goal-reaching reward $r_{g}^{T}$, a success bonus $r_{s}^{T}$, and a negative affordance penalty $r_{n}^{T}$. Each component is detailed below.
The grasp reward $r_{d}^{T}$ encourages the hand to approach and stay close to the object. It is defined as a penalty proportional to the distance between the hand and the object’s center:
$r_{d}^{T}=\lambda_{d}^{T}\|p_{\text{dex}}-p_{\text{obj}}\|,$ (9)
where $p_{\text{dex}}$ and $p_{obj}$ are the Cartesian position of the dexterous hand and the object, respectively. We set $\lambda_{d}^{T}$ to -1.
The goal reward $r_{g}^{T}$ guides the object towards the target goal. It is structured as a penalty on the distance between the object and the goal:
$r_{g}^{T}=\lambda_{g}^{T}\|p_{\text{obj}}-p_{\text{goal}}\|,$ (10)
where $p_{\text{goal}}$ is the position of the target goal. We set $\lambda_{g}^{T}$ to -1.
A bonus is provided upon task completion. This is triggered when the object enters a small threshold radius around the goal:
$r_{s}^{T}=\lambda_{s}^{T}\mathbb{I}(\|p_{\text{obj}}-p_{\text{goal}}\|<\alpha_{s}),$ (11)
where $\mathbb{I}(\cdot)$ is the indicator function, which evaluates to 1 if the condition is true and 0 otherwise. $\alpha_{s}$ is set to 0.05. We set $\lambda_{s}^{T}$ to 1.
The negative affordance reward $r_{n}^{T}$ penalizes the dexterous hand fingertips to approach the predicted negative affordance:
$r_{n}^{T}=\lambda_{n}^{T}\sum_{f\in\mathcal{K}_{c}}\mathbb{I}\left(\min_{p_{n}\in\mathcal{P}_{n}}\|p_{\text{tip}}^{f}-p_{n}\|<\alpha_{n}\right),$ (12)
where $\alpha_{n}$ is set to 0.03, $\mathcal{K}_{c}$ is the set of the hand’s fingertips, $p_{\text{tip}}^{f}$ is the position of fingertip $f$, and $\mathcal{P}_{n}$ is the set of points representing the negative affordance. We set $\lambda_{n}^{T}=-10$. Ablation studies are conducted on these weights to evaluate their contribution. The results are presented in Hyperparameters part.
Table 5: Ablation Study on UniDexGrasp dataset (xu2023unidexgrasp) in Seen Object in state-based setting.
Configuration $Succ\uparrow$ $HLS\uparrow$ $AS\downarrow$
$\lambda_{d}^{T}=-0.5$ 84.2 7.2 5
$\lambda_{d}^{T}=-1$ 89.2 8.6 4
$\lambda_{d}^{T}=-2$ 88.4 8.4 4
$\lambda_{g}^{T}=-0.5$ 87.9 8.1 4
$\lambda_{g}^{T}=-1$ 89.2 8.6 4
$\lambda_{g}^{T}=-2$ 88.9 8.3 6
$\lambda_{s}^{T}=0.5$ 83.1 7.0 6
$\lambda_{s}^{T}=1$ 89.2 8.6 4
$\lambda_{s}^{T}=2$ 89.0 8.5 5
$\lambda_{n}^{T}=-5$ 89.1 8.5 16
$\lambda_{n}^{T}=-10$ 89.2 8.6 4
$\lambda_{n}^{T}=-20$ 87.1 8.0 0
Human-likeness Score ($HLS$). The Human-likeness Score (HLS) is designed to quantify the anthropomorphic quality of the generated grasping motion. We leverage the advanced multi-modal capabilities of Gemini 2.5 Pro (comanici2025gemini) to serve as an expert evaluator.
For each grasp, a video sequence of the complete grasp execution is provided as input to the model. The model is then prompted with a specific set of instructions to assess the motion based on key kinematic criteria. The exact prompt provided to the model is as follows:
Expert Role: You are an expert in hand kinematics evaluation.
Task: Evaluate the similarity between the robotic hand motion in the simulation video and that of a real human hand, based on the following three criteria:
  • Motion trajectory
  • Velocity smoothness
  • Joint coordination
Output Format: Return only a valid JSON, following this format:
{ "score": <1-10> }.
VLM-based Negative Affordance Identification VLMs struggle to interpret non-textured 3D meshes, as these models primarily rely on rich visual cues learned from images. To bridge this gap, we first apply procedural texturing to the raw meshes using Tex-Painter (zhang2024texpainter), which generates semantically plausible textures based on geometric analysis. Next, we render the textured object from six cardinal directions to create a multi-view image set, $I$, as a holistic visual representation. While this may not capture all concavities in highly complex objects, we found it provides a sufficient basis for affordance prediction for the objects in our benchmark datasets.
We then query GPT-4V (achiam2023gpt) with the multi-view images to elicit a description of the object’s negative affordances. The exact prompt is as follows:
Expert Role: You are an expert in identifying which part of a physical object should not be touched, especially in robotic grasping tasks.
Task: You will be given 6-view images of an object. Your task is:
  • Identify one non-touchable part.
  • Infer the object’s identity from the image name.
Output Format: Return one sentence only, following this format:
"This is a _____. _____ of ______ should not be touched."
Points selected in Negative Affordance-aware Segmentation The core idea is to systematically prompt SAM with a dense, regular grid of points covering the entire image. This process ensures that objects of various sizes and locations are likely to be ”hit” by at least one point prompt, triggering SAM to segment them. The procedure consists of the following steps:
  1. 1.
    Grid Definition: We define a regular grid of points to be overlaid on the input image. The density of this grid is controlled by a single hyperparameter, $g=16$, which represents the number of points along each of the image’s dimensions. This results in a total of $g^{2}$ keypoints.
  2. 2.
    Keypoint Generation: For an input image $I$ with width $W$ and height $H$, the set of grid points $G$ is generated. The coordinates $(x_{i},y_{j})$ for each point are calculated by uniformly spacing them across the image dimensions. The formula for a point at grid position $(i,j)$ is:
    $G=\left\{(x_{i},y_{j})\mid x_{i}=\frac{i}{g+1}\cdot W,\quad y_{j}=\frac{j}{g+1}\cdot H\right\}$ (13)
    for all $i,j\in\{1,2,\dots,g\}$. Using $g+1$ as the denominator ensures that the points are placed within the image boundaries and not on the absolute edges.
  3. 3.
    Prompting SAM: Each of the $g^{2}$ points in the set $G$ is used as an individual positive point prompt for SAM. This yields a large collection of raw segmentation masks, $\mathcal{M}_{\text{raw}}=\text{SAM}(I,G)$, where many masks may be redundant or highly overlapping.
  4. 4.
    Filtering with NMS: To produce a refined set of unique proposals, we apply Non-Maximum Suppression (NMS) to the raw masks. The NMS algorithm filters out duplicate masks based on the Intersection over Union (IoU) of their corresponding bounding boxes, retaining only the most confident and distinct object proposals.
Hyperparameters. As shown in Tab. 5, we conduct a series of ablation studies to validate our design choices and analyze the model’s sensitivity to key hyperparameters. Our investigation focuses on the weights of our proposed reward function. The goal is to find a robust set of parameters that balances efficient learning with task performance.
The results indicate that our chosen default configuration is robust and well-justified. For the grasp reward weight $\lambda^{T}_{d}$ and the goal reward weight $\lambda_{g}^{T}$, we observe an optimal value of -1.0. Deviating from this value, by either decreasing the penalty to -0.5 or increasing it to -2.0, leads to a degradation in both $Succ$ and $HLS$. This suggests a critical balance: a penalty that is too weak fails to prevent undesirable actions, while an overly strong penalty can cause the policy to focus myopically on avoiding penalties rather than achieving a stable grasp, ultimately degrading task success. Similarly, for the bonus reward weight $\lambda^{T}_{s}$, the optimal value is 1.0. A lower weight of 0.5 fails to sufficiently incentivize the target behavior, resulting in a significant drop in performance. Conversely, increasing the weight to 2.0 offers no additional benefit and slightly hinders performance, indicating that an excessive bonus can also be detrimental. Similarly, the negative affordance weight $\lambda_{n}^{T}$ requires careful tuning. An overly strong penalty can make the agent too conservative, causing it to avoid the target object altogether in its effort to steer clear of negative affordances. Conversely, an insufficient penalty fails to effectively deter the agent from frequently approaching these undesirable regions.
10 Future Experiments
We chose simulation to enable large-scale generalization experiments. The policy is vision-based and avoids privileged information, designing it for sim-to-real transfer, which is a key future goal.